OpenStudy (anonymous):

Medal&Fan. Solve the equation Quantity x plus 1 all over 3 = 2 for x. Please show your work and solution in the space provided.

OpenStudy (nightwielder):

\[\frac{ x+1 }{ 3 }=2\]

OpenStudy (nightwielder):

Multiply both sides by denominator: \[x+1=6\]

OpenStudy (nightwielder):

Subtract both sides by 1: \[x=5\]

OpenStudy (nightwielder):

Final Answer: \[x=5\]